# Distilling 3D distinctive local descriptors for 6D pose estimation

### v2

@@ V-A Experimental setup
We validate dGeDi with three experiments: (i) a direct comparison against GeDi [2], (ii) a detailed analysis of dGeDi’s key components, and (iii) a comparison with state-of-the-art zero-shot pose estimators by integrating dGeDi in FreeZe [3]. During preprocessing, we randomly sample 4k points from $\mathcal{P}^{Q}$ and $\mathcal{P}^{T}$, and apply statistical outlier removal [57] to $\mathcal{P}^{T}$ to reduce depth noise. Both point clouds are normalized by $\mathcal{P}^{Q}$’s diameter to promote scale invariance. In (i) and (ii), we perform feature extraction using only the geometric encoder, while in (iii) we integrate dGeDi with FreeZe’s vision encoder. In (i) and (iii), we use both ground-truth and SAM-6D [29] segmentation masks to ensure fairness and compliance with the zero-shot assumption. In (ii), we perform object localization using ground-truth segmentation masks to focus the analysis on feature quality without localization errors. GeDi extracts 32-dimensional point-level features from local neighborhoods covering 30% of $Q$’s diameter. After feature extraction, we perform feature matching using the Open3D’s RANSAC and ICP implementations. We run 100k iterations for RANSAC and 1k iterations for ICP. The student model is trained using eight NVIDIA A100 GPUs, while evaluation is performed on a single NVIDIA A40 GPU. Note that the peak memory usage of dGeDi when processing a point cloud with 5k points is only 717MB, making it suitable for deployment on edge devices with limited resources.
Training dataset. We train dGeDi on the MegaPose-ShapeNetCore [26] dataset, which contains 1 million synthetic images generated by rendering various 3D objects against random background scenes. The dataset provides ground-truth segmentation masks, which we use to localize target objects.
@@ V-B Quantitative results
GeDi vs. dGeDi. Tab. I compares GeDi [2] (teacher, Rows 1-12) and dGeDi (student, Rows 13-24) on the BOP Benchmark using different types of segmentation masks. When using ground-truth segmentation masks, dGeDi significantly outperforms GeDi by +4.1 RON, +11.8 FMR, and +2.4 AR, while being 170 times faster (Row 25). dGeDi outperforms GeDi in RON and FMR scores across all datasets but shows a slight drop in AR on TUD-L[60] and YCB-V[62]. This difference arises because RON and FMR directly assess feature quality, whereas AR is computed post registration, where RANSAC’s robustness to outliers can obscure the true contribution of the features. On T-LESS [59], dGeDi achieves a substantial +14.1 AR improvement over the teacher. We attribute this significant gain to the specific type of 3D models: unlike other datasets, T-LESS provides CAD models of industrial objects with fine-grained geometric details and sharp edges. As shown in Fig. 3, dGeDi excels in this setting, producing features that are smoother that the teacher’s and better capture the underlying geometry. When using SAM-6D zero-shot segmentation masks, dGeDi outperforms the teacher by +2.7 RON and +9.4 FMR, but achieves a -1.1 AR (Row 26). This setting is more challenging than the former because zero-shot segmentation masks may have inaccuracies along object contours, can miss some parts of the object, and can include parts of other objects or background, while during distillation we use ground-truth segmentation masks. Despite this, dGeDi proves to be robust to the domain gap, achieving a superior feature quality compared to the teacher (better RON and FMR) and a comparable AR. Beyond improved accuracy, dGeDi also achieves superior computational efficiency, extracting features at more than 50 frames per second. Given that each image contains an average of five target objects, this implies that dGeDi can process approximately 260 point clouds per second. Processing times (rightmost column) depends on the number of object instances per image, ranging from 3ms for a single instance in TUD-L to 34ms for dozens of instances in IC-BIN [61]. In summary, dGeDi delivers high accuracy with low latency, making it suitable for time-critical robotics applications.
@@
23 Avg 7.1 36.3 62.2 0.019
24 5.6 27.9 51.7
25 $\Delta$ +4.1 +11.8 +2.4 173$\times$
26 +2.7 +9.4 -1.1
@@
Ablation study. Tab. II analyses the impact of three key components in dGeDi: supervision type (Col. 1), loss function design (Cols. 2-5), and data augmentation strategy (Col. 6-8). Row 1 is our baseline, which uses supervision only from query objects, Eq. 1 as loss, and no data augmentation. Row 9 is dGeDi’s final configuration, while Rows 2-8 explore alternative designs. We evaluate each variant on the BOP Benchmark and report its improvement over the baseline in the rightmost column. We draw three main insights for these experiments. First, the learning via correspondence paradigm introduced in Sec. IV-A improves performance considerably (+5.2 AR, Row 5) by introducing supervision from partial and noisy target objects with the same memory requirements. Second, the proposed loss (Eq. 1) outperforms both traditional designs (+3.9 AR against MSE in Row 2, +3.5 AR against Huber in Row 3) and Eq. 2 (+8.7 AR, Row 4). This indicates that effective distillation of frozen GeDi features requires a careful loss design, and that reducing the contribution of points with a large regression error is beneficial. Third, data augmentation promote generalization (Row 9 achieves +0.4 AR over Row 5). These ablation experiments confirm that integrating correspondence-based supervision, a tailored loss, and data augmentation consistently enhances performance.
@@
dGeDi and FreeZe integration. Tab. III assesses the 6D pose estimation performance obtained when integrating dGeDi in FreeZe. We compare against state-of-the-art zero-shot 6D pose estimators on the same datasets as Tab. I, evaluating AR and run-time. We compare the performance of dGeDi with SAM-6D [29], FoundationPose [30], and the original version of FreeZe, which are the state-of-the-art zero-shot methods. Rows 1-3 present publicly available results from the BOP Benchmark leaderboard [63]. Rows 4-5 present the reproduced results of FreeZe, both with and without symmetry-aware pose refinement (SAR). Rows 6-7 show performance when replacing FreeZe’s geometric encoder GeDi [2] with dGeDi. To ensure a fair comparison, all methods use the SAM-6D zero-shot segmentor for object localization except for Rows 8-9 where ground-truth segmentation masks are used. For experiments using ground-truth segmentation masks, dGeDi achieves the same AR of FreeZe, while reducing the inference time by 2.7 seconds (Row 9). When using SAM-6D segmentation masks, dGeDi reaches 70.4 AR compared to FreeZe’s 71.5, yet it runs 5.2 seconds faster. This run-time improvement is mainly due to dGeDi ability to process multiple target objects per image in parallel, unlike FreeZe, which processes them serially. In contrast, the slight performance drop is mainly due to noise in the predicted segmentation masks. Since dGeDi is trained with ground-truth, noise-free segmentation masks, the introduction of noisy SAM-6D segmentation masks creates a domain gap that leads to suboptimal pose estimation, although dGeDi still maintains better feature quality and faster processing.

### v3

@@ V-A Experimental setup
We validate dGeDi with three experiments: (i) a direct comparison against GeDi [2], (ii) a detailed analysis of dGeDi’s key components, and (iii) a comparison with state-of-the-art zero-shot pose estimators by integrating dGeDi into FreeZe [3]. During preprocessing, we randomly sample 4k points from $\mathcal{P}^{Q}$ and $\mathcal{P}^{T}$, and apply statistical outlier removal [57] to $\mathcal{P}^{T}$ to reduce depth noise. Both point clouds are normalized by $\mathcal{P}^{Q}$’s diameter to promote scale invariance. In (i) and (ii), we perform feature extraction using only the geometric encoder, while in (iii) we integrate dGeDi with FreeZe’s vision encoder. In (i) and (iii), we use both ground-truth and SAM-6D [29] segmentation masks to ensure fairness and compliance with the zero-shot assumption. In (ii), we perform object localization using ground-truth segmentation masks to focus the analysis on feature quality without localization errors. GeDi extracts 32-dimensional point-level features from local neighborhoods covering 30% of $Q$’s diameter. After feature extraction, we perform feature matching using the Open3D’s RANSAC and ICP implementations. We run 100k iterations for RANSAC and 1k iterations for ICP. The student model is trained using eight NVIDIA A100 GPUs, while evaluation is performed on a single NVIDIA A40 GPU. Note that the peak memory usage of dGeDi when processing a point cloud with 5k points is only 717MB, making it suitable for deployment on edge devices with limited resources. The student architecture is adapted from PTV3 [18], using the same model settings but modifying the decoder output channels from 64 to 32 to match the teacher feature dimensionality. The model is trained using the AdamW optimizer with a cosine learning rate scheduler.
Training dataset. We train dGeDi on the MegaPose-ShapeNetCore [26] dataset, which contains one million synthetic images generated by rendering various 3D objects against random background scenes. The dataset provides ground-truth segmentation masks, which we use to localize target objects.
@@ V-B Quantitative results
23 Avg 7.1 36.3 66.8 0.019
24 5.6 27.9 56.8
[…]
26 +2.7 +9.4 +1.7
GeDi vs. dGeDi. Tab. I compares GeDi [2] (teacher, Rows 1-12) and dGeDi (student, Rows 13-24) on the BOP Benchmark using different types of segmentation masks. When using ground-truth segmentation masks, dGeDi significantly outperforms GeDi by +4.1 RON, +11.8 FMR, and +3.9 AR, while being 170 times faster (Row 25). dGeDi outperforms GeDi in RON and FMR scores across all datasets but shows a slight drop in AR on TUD-L[60] and YCB-V[62]. This difference arises because RON and FMR directly assess feature quality, whereas AR is computed post registration, where RANSAC’s robustness to outliers can obscure the true contribution of the features. On T-LESS [59], dGeDi achieves a substantial +22.9 AR improvement over the teacher. We attribute this significant gain to the specific type of 3D models: unlike other datasets, T-LESS provides CAD models of industrial objects with fine-grained geometric details and sharp edges. As shown in Fig. 3, dGeDi excels in this setting, producing features that are smoother that the teacher’s and better capture the underlying geometry. When using SAM-6D zero-shot segmentation masks, dGeDi outperforms the teacher for all three metrics by +2.7 RON, +9.4 FMR and +1.7 AR (Row 26). This setting is more challenging than the former because zero-shot segmentation masks may have inaccuracies along object contours, can miss some parts of the object, and can include parts of other objects or background, while during distillation we use ground-truth segmentation masks. Despite this, dGeDi proves to be robust to the domain gap, achieving a superior feature quality compared to the teacher (better RON and FMR) and a superior AR. Beyond improved accuracy, dGeDi also achieves superior computational efficiency, extracting features at more than 50 frames per second. Given that each image contains an average of five target objects, this implies that dGeDi can process approximately 260 point clouds per second. Processing times (rightmost column) depends on the number of object instances per image, ranging from 3ms for a single instance in TUD-L to 34ms for many instances in IC-BIN [61]. In summary, dGeDi combines high accuracy and low latency, making it suitable for robotics applications.
@@
Ablation study. Tab. II analyses the impact of three key components in dGeDi: supervision type (Col. 1), loss function design (Cols. 2-5), and data augmentation strategy (Col. 6-8). Row 1 is our baseline, which uses supervision only from query objects, Eq. 1 as loss, and no data augmentation. Rows 2-8 explore alternative designs, while Row 9 is dGeDi’s final configuration. We evaluate each variant on the BOP Benchmark and report its improvement over the baseline in the rightmost column. We draw three main insights for these experiments. First, the learning via correspondence paradigm introduced in Sec. IV-A improves performance considerably (+4.7 AR, Row 5) by introducing supervision from partial and noisy target objects with the same memory requirements. Second, the proposed loss (Eq. 1) outperforms both traditional designs (+3.4 AR against MSE in Row 2, +3.1 AR against Huber in Row 3) and Eq. 2 (+8.5 AR, Row 4). This indicates that effective distillation of frozen GeDi features requires a careful loss design, and that reducing the contribution of points with a large regression error is beneficial. Third, data augmentation promote generalization (Row 9 achieves +1.3 AR over Row 5). These ablation experiments confirm that integrating correspondence-based supervision, a tailored loss, and data augmentation consistently enhances performance.
@@
dGeDi and FreeZe integration. Tab. III assesses the 6D pose estimation performance obtained when integrating dGeDi in FreeZe. We compare against state-of-the-art zero-shot 6D pose estimators on the same datasets as Tab. I, evaluating AR and run-time. We compare the performance of dGeDi with SAM-6D [29], FoundationPose [30], and the original version of FreeZe, which are the state-of-the-art zero-shot methods. Rows 1-3 present publicly available results from the BOP Benchmark leaderboard [63]. Rows 4-5 present the reproduced results of FreeZe, both with and without symmetry-aware pose refinement (SAR). Rows 6-7 show performance when replacing FreeZe’s geometric encoder GeDi [2] with dGeDi. To ensure a fair comparison, all methods use the SAM-6D zero-shot segmentor for object localization except for Rows 8-9 where ground-truth segmentation masks are used. For experiments using ground-truth segmentation masks, dGeDi achieves the same AR of FreeZe, while reducing the inference time by 2.7 seconds (Row 9). When using SAM-6D segmentation masks, dGeDi reaches 70.4 AR compared to FreeZe’s 71.5, yet it runs 5.2 seconds faster. In contrast, the slight performance drop is mainly due to noise in the predicted segmentation masks. Since dGeDi is trained with ground-truth, noise-free segmentation masks, the introduction of noisy SAM-6D segmentation masks creates a domain gap that leads to suboptimal pose estimation, although dGeDi still maintains better feature quality and faster processing.